# On the Reachability of 3–Dimensional Paths with a Prescribed Curvature Bound

### v2

@@ 4 Conclusion
In this paper, we have presented the reachability analysis of three–dimensional curves with a prescribed curvature bound. We introduced the set of time–optimal control problems to show the internal connection between time–optimality and reachability through Pontryagin Maximum Principle. Leveraging this relationship, we provided an implicit description of the boundary of the reachability set, showing that any boundary point can be reached by curves belonging to the classes CSC, CCC, their subsegments, or H. For the position reachability set, we proved that any boundary point can be reached by planar curves belonging to the classes CS, CC, or their subsegments. Moreover, we proved that the reachability set in this case is a solid of revolution of its two–dimensional counterpart, the Dubins car, thereby enabling complete determination. These developments extend the existing literature on Dubins car into dimension 3. As the foundational studies on the reachability set of Dubins car did, such generalization into 3D curves is anticipated to enable various advancements in guidance and path planning methods. {ack} We would like to extend our gratitude to Nearthlab for insights and expertise during the review process of our paper.

### v3

@@ 4 Conclusion
In this paper, we have presented the reachability analysis of three-dimensional curves with a prescribed curvature bound. Based on Pontryagin Maximum Principle and existing results on Markov-Dubins problem in dimension 3, we provided an implicit description of the boundary of the reachability set that any boundary point can be reached by curves belonging to the classes CSC, CCC, their subsegments, or H. We further proved that the position reachability set is a solid of revolution of its two-dimensional counterpart, the Dubins car, thereby enabling complete determination. These developments extend the existing literature on Dubins car into dimension 3. {ack} We would like to extend our gratitude to Nearthlab for supports for this paper, and the anonymous reviewers for their insightful comments.